A) During February, $84,200 was paid to creditors on account, and purchases on account were $107,780. Assuming that the February 28 balance of Accounts Payable was $36,210, determine the account balance on February 1. $

b. On October 1, the accounts receivable account balance was $38,900. During October, $338,400 was collected from customers on account. Assuming that the October 31 balance was $44,700, determine the fees billed to customers on account during October. $

c. On April 1, the cash account balance was $16,640. During April, cash receipts totaled $252,930 and the April 30 balance was $11,980. Determine the cash payments made during April. $

The following errors took place in journalizing and posting transactions:

  1. Insurance of $7,200 paid for the current year was recorded as a debit to Insurance Expense and a credit to Prepaid Insurance.
  2. A withdrawal of $22,000 by Brian Phillips, owner of the business, was recorded as a debit to Wages Expense and a credit to Cash.

Journalize the entries to correct the errors. For part a, first reverse the original entry and then make the correct entry.

a.
b.
Answer #1

Part 1 A

Opening balance + Purchases- Cash paid= Ending balance

Therefore,

Opening balance = Ending balance+ Cash paid-Purchases= $36210+$84200-$107780 = $12630

Account balance on February 1= $12630

Part 1 B

Beginning Balance + Credit Sales - Collections = Ending Balance

Therefore,

Credit sales = Ending Balance - Beginning Balance + Collections = 44700-38900+338400 = $344200

Fees billed to customers on account during October = $344200

Part 1 C

Cash payments = Beginning balance + cash receipts – ending balance = 16640+252930-11980 = $257590

Part 2

No.

Accounts title and explanation

Debit

Credit

a

Prepaid insurance

14400

Insurance expense

7200

cash

7200

b

Drawings

22000

Wages expense

22000